OpenStudy (aonz):

Suppose that the cost of producing x items per hour is given by C(x) where C(x) = x^2+10, and the number of items sold per hour at a price of p dollars per item is x = 16 − p. (a) Find in terms of x the revenue gained from the sales.

OpenStudy (anonymous):

Okay, so if I tell you that revenue is \(px\), does it make sense to you?

OpenStudy (anonymous):

\(p\) is the price, which is also the revenue per item. \(x\) is the number of items

OpenStudy (aonz):

yea

OpenStudy (anonymous):

\[ x = 16 − p \]Multiplying both sides by \(x\) get you:\[ x^2 = 16x-px \]

OpenStudy (anonymous):

And solving for \(px\):\[ x^2 = 16x - px \\ px + x^2 = 16x \\ px = 16x-x^2 \]

OpenStudy (anonymous):

Does that make sense to you?

OpenStudy (aonz):

yep so far

OpenStudy (anonymous):

\[ R(x) = px \implies R(x) = 16x-x^2 \]

OpenStudy (anonymous):

The answer is already there

OpenStudy (aonz):

oh so the question didnt need x^2 + 10

OpenStudy (anonymous):

Not part (a)

OpenStudy (aonz):

mind if i post the full question?

OpenStudy (anonymous):

Go ahead

OpenStudy (aonz):

(b) Hence show that the profit achieved per hour is given by −2x^2 +16x−10. (c) Find the number of items that should be produced each hour in order to maximise the profit. (d) Find the maximum profit. I can do part (d) so dw about that part

OpenStudy (anonymous):

Okay, well total profit (or just profit) is just total cost minus total revenue

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

What is the formula for profit?

OpenStudy (aonz):

px?

OpenStudy (anonymous):

No, that was the formula for revenue.

OpenStudy (aonz):

px - c(x)

OpenStudy (anonymous):

Okay, so then substitute in what you solved, and see what it simplifies to...

OpenStudy (aonz):

yea same as that part(b) gave us ^^

OpenStudy (anonymous):

Okay so now for c

OpenStudy (anonymous):

What is \(x\) such that \(−2x^2 +16x−10\) is maximized?

OpenStudy (aonz):

hmm?

OpenStudy (aonz):

4?

OpenStudy (anonymous):

LOL, how did you get that?

OpenStudy (aonz):

x = -b/2a

OpenStudy (anonymous):

Eh? Is that how you get the max? I've never seen that... Interesting.

OpenStudy (aonz):

x = 4 so max = 22

OpenStudy (anonymous):

Let me think aloud here.... \[ ax^2+bx+c=f(x) \\ f'(x) = 0 \implies 2ax+b = 0 \implies x = -\frac{b}{2a} \]Oh wow, you're right, good job.

OpenStudy (aonz):

ok anyways thanks for your help again ^^

OpenStudy (anonymous):

Interesting formula though.

OpenStudy (aonz):

In Australia it is a formula you need to memorise.